OpenStudy (anonymous):

**MEDAL + FAN** One phone plan charges a $20 monthly fee and $0.08 per minute on every phone call made. Another phone plan charges a $12 monthly fee and $0.12 per minute for each call. After how many minutes would the charge be the same for both plans? A. 60 minutes B. 90 minutes C. 120 minutes D. 200 minutes <-- Formula?

OpenStudy (gorv):

cost=20+x*0.08 another phone cost=12+x*0.12

OpenStudy (gorv):

x=minute of call both to be equal 20+x*0.08=12+x*0.12

OpenStudy (gorv):

20+0.08x=12+0.12x subtract 12 from both side 8+.08x=0.12x subtact 0.08x from both side 8=0.04x divide by 0.04 both side

OpenStudy (gorv):

x=8/0.04=200

OpenStudy (anonymous):

@gorv Thanks! I'll fan you once my Openstudy client stops bugging out.

OpenStudy (gorv):

welcome